Volcán Barva (10.01.2012 to 12.31.2012)

You are here

PDF versionPDF version
Year: 
2012
Quarter: 
October - December
Communications activities: 

There were not communications activities this period.

Unusual Events at the TEAM Site: 

There were not unusual events this period.

New Species: 

There were not new species this period.

Protocol Activities: 

During the last 3 months, climate data were collected every month.

We started the vegetation sampling period on December for the plot locate at 200masl. Calibration data was also done.

Protocol Problems: 

Things are running in the Climate station and data have been downloaded, but there were two problems:The first is that the temperature data from the second HMP45 sensor is bad. This is the same sensor were we lost relative humidity data earlier. Second is that night time battery voltages dropped below 9V (since 20 Nov or so) which caused the logger to stop. This means that night time data is missing for a lot of the night during the last month. During daytime it does not seem a big problem. It was one of the smaller batteries that was defect and was drawing current from the other two batteries (it was no longer functioning as battery, but rather as load). I disconnected the bad battery was disconnected and the system should keep 11V at night now.

We have schedule in the first months of 2013 year, the instalation of a new big battery in order to improve the climate station performance.

Schedule Problems: 

There were not schedule problems.

Logistical Problems: 

There were not logistical problems during this period.

Suggestions: 

There is not suggestions this time.